# STOPA: A Dataset of Systematic VariaTion Of DeePfake Audio for Open-Set Source Tracing and Attribution

### v2

@@ 1 Introduction
A key difference of STOPA compared to existing datasets and most of the published methodology literature relies on the evaluation protocols design, intended to promote development of source tracing methods for ‘open-world’ settings. To elaborate, the majority of the source-tracing literature approaches the task as a multiclass task  [4, 5, 9, 10, 11] assuming known attacks during training and handling unknown attacks via an additional fallback class with a dedicated out-of-distribution classifier. Whereas this treatment leads to class priors dependencies in performance metrics, extra classifiers further complicate system design and evaluation.
Inspired by the established principles of the speaker detection evaluation benchmarks coordinated by NIST [12], we frame source tracing as an open-world detection task that unifies source tracing and ’detecting the unknown’ under a simple and coherent evaluation framework. A test instance is evaluated against a hypothesized spoofing attack, vocoder, or acoustic model. The task is to answer whether or not the test instance originates from the hypothesized source generator (null hypothesis) or not (alternative hypothesis). Similar to NIST SREs, designed to support ’continual growing’ of a speaker database, our protocol forbids the use of other attacks for training or testing purposes: a decision must solely be made by comparing the hypothesized model against a single test instance. This mindset intends to promote the design of deepfake profiling approaches where the database of deepfake ’signatures’ (similar to databases of computer viruses or malware) is allowed to grow dynamically during the lifespan of the profiling system. A ’fixed’ dataset of deepfake generators, addressed as a multiclass task, requires re-training the profiling system whenever a new deepfake generator is encountered.
STOPA’s main contributions include: (1) a scalable evaluation protocol enabling open-world source tracing by integrating new attack signatures without retraining, (2) systematic variation in spoofed speech with diverse acoustic models, vocoders, and hyperparameters, and (3) extensive metadata for fine-grained analysis and benchmarking of synthetic speech.
@@ 3.2 Post-processing
For each waveform $\mathbf{x}$, we removed DC bias and normalized amplitude before detecting non-speech segments. Speech onset $s^{*}$ and offset $e^{*}$ were determined based on voice activity detection (VAD) activation or root mean square energy (RMSE) exceeding a threshold $\theta_{\text{rmse}}$. Leading and trailing silence were identified by counting consecutive non-speech frames, with stable RMSE over $W$ frames signalling the end of the speech. Empirical thresholds ($\theta_{\text{rmse}}=0.05,W=18$) were optimized to remove non-speech while preserving intelligibility. We evaluated intelligibility using Word Error Rate (WER) with Whisper-large-v3 [41] and JiWER [42]. Samples with WER $>$ 100 (1%) were removed from train/dev sets while preserving speaker balance but retained in eval, as attribution models should also be able to handle degraded inputs.
@@ 4 Demonstration of STOPA Use
We address attack type, acoustic model, and vocoder model classification using three ’trace embedding’ extractors: AASIST CM [45] countermeasure (trained on ASVspoof2019 LA for binary spoof detection), and ResNet-34 [46] and AASIST STOPA, both trained on the STOPA training partition for attribution. ResNet-34 and AASIST STOPA are trained for 100 epochs. We emphasize that none of these models have been exposed to the attacks used in the source tracing reference and trial sets, as the STOPA dataset is explicitly designed to be attack-disjoint. The ‘zero-shot’ mindset (the ability to add new attacks without needing to retrain the embedding extractor) is a key design factor of STOPA protocols.
The actual trace embeddings for each attack are extracted using the predefined models. Following the simplest known classifier from speaker detection literature, we form a model for each attack by averaging the training embeddings. Test utterances are then compared to each reference using cosine similarity. A positive trial is where the ground-truth label (attack, vocoder, or acoustic model ID) in the model and the test match; otherwise, it is a negative trial. The negative set is further split into two subsets: (1) a different known attack and (2) an unknown attack, allowing for EER computation in both scenarios. The latter represents ’the unknown unknown’.
Table 4 shows error rates for detecting attack type (ATK), acoustic model (AM), and vocoder model (VM). The very high EERs reflects the difficulty of our open-world task setting. The best performance was obtained with the ASVspoof2019-trained AASIST CM. Both AASIST STOPA and the ResNet-34 models solely trained on STOPA produce, essentially, a chance rate. Since ASVspoof 2019 contains a larger number of training attacks, this strongly suggests that the more limited set of 3 STOPA attacks is not sufficient for capturing ’general variability’ of attack-specific traits. Furthermore, using common vs. non-common trace embedding utterances has no impact on final performance.
@@
It might be useful to reflect these results on those typically obtained from speaker detection studies. Modern speaker embedding extractors, such as ECAPA-TDNN, are typically trained using thousands of speakers (classes). This leads to a discriminative ’speaker trait’ space that transfers easily to classify previously unseen speakers. In contrast, we used only a handful of attacks (3 in STOPA) in training.
@@ 5 Conclusions
We introduced STOPA, a novel dataset for synthetic speech source tracing, featuring 13 distinct attacks with systematic variations in acoustic and vocoder models. Unlike existing datasets, STOPA is designed to address ’open-world’ source tracing. Inspired by NIST speaker detection benchmarks, our evaluation protocols unify source tracing and unknown attack detection under a scalable framework. By prohibiting the use of other attack data for training or testing, we encourage the development of dynamic ’attack signature databases’ that evolve over time, similar to antivirus systems.

### v3

@@ 1 Introduction
A key difference of STOPA compared to existing datasets and most of the published methodology literature relies on the evaluation protocols design, intended to promote development of source tracing methods for `open-world' settings. To elaborate, the majority of the source-tracing literature approaches the task as a multiclass task  [4, 5, 9, 10, 11] assuming known attacks during training and handling unknown attacks via an additional fallback class with a dedicated out-of-distribution classifier. Whereas this treatment leads to class priors dependencies in performance metrics, extra classifiers further complicate system design and evaluation.
Inspired by the established principles of the speaker detection evaluation benchmarks coordinated by NIST [12], we frame source tracing as an open-world detection task that unifies source tracing and 'detecting the unknown' under a simple and coherent evaluation framework. A test instance is evaluated against a hypothesized spoofing attack, vocoder, or acoustic model. The task is to answer whether or not the test instance originates from the hypothesized source generator (null hypothesis) or not (alternative hypothesis). Similar to NIST SREs, designed to support 'continual growing' of a speaker database, our protocol forbids the use of other attacks for training or testing purposes: a decision must solely be made by comparing the hypothesized model against a single test instance. This mindset intends to promote the design of deepfake profiling approaches where the database of deepfake 'signatures' (similar to databases of computer viruses or malware) is allowed to grow dynamically during the lifespan of the profiling system. A 'fixed' dataset of deepfake generators, addressed as a multiclass task, requires re-training the profiling system whenever a new deepfake generator is encountered.
STOPA's main contributions include: (1) a scalable evaluation protocol enabling open-world source tracing by integrating new attack signatures without retraining, (2) systematic variation in spoofed speech with diverse acoustic models, vocoders, and hyperparameters, and (3) extensive metadata for fine-grained analysis and benchmarking of synthetic speech.
@@ 3.2 Post-processing
For each waveform $\mathbf{x}$, we removed DC bias and normalized amplitude before detecting non-speech segments. Speech onset $s^{*}$ and offset $e^{*}$ were determined based on voice activity detection (VAD) activation or root mean square energy (RMSE) exceeding a threshold $\theta_{\text{rmse}}$. Leading and trailing silence were identified by counting consecutive non-speech frames, with stable RMSE over $W$ frames signalling the end of the speech. Empirical thresholds ($\theta_{\text{rmse}}=0.05,W=18$) were optimized to remove non-speech while preserving intelligibility. We evaluated intelligibility using Word Error Rate (WER) with Whisper-large-v3 [41] and JiWER [42]. Samples with WER >> 100 (1%) were removed from train/dev sets while preserving speaker balance but retained in eval, as attribution models should also be able to handle degraded inputs.
@@ 4 Demonstration of STOPA Use
We address attack type, acoustic model, and vocoder model classification using three 'trace embedding' extractors: AASIST CM [45] countermeasure (trained on ASVspoof2019 LA for binary spoof detection), and ResNet-34 [46] and AASIST STOPA, both trained on the STOPA training partition for attribution. ResNet-34 and AASIST STOPA are trained for 100 epochs. We emphasize that none of these models have been exposed to the attacks used in the source tracing reference and trial sets, as the STOPA dataset is explicitly designed to be attack-disjoint. The `zero-shot' mindset (the ability to add new attacks without needing to retrain the embedding extractor) is a key design factor of STOPA protocols.
The actual trace embeddings for each attack are extracted using the predefined models. Following the simplest known classifier from speaker detection literature, we form a model for each attack by averaging the training embeddings. Test utterances are then compared to each reference using cosine similarity. A positive trial is where the ground-truth label (attack, vocoder, or acoustic model ID) in the model and the test match; otherwise, it is a negative trial. The negative set is further split into two subsets: (1) a different known attack and (2) an unknown attack, allowing for EER computation in both scenarios. The latter represents 'the unknown unknown'.
Table 4 shows error rates for detecting attack type (ATK), acoustic model (AM), and vocoder model (VM). The very high EERs reflects the difficulty of our open-world task setting. The best performance was obtained with the ASVspoof2019-trained AASIST CM. Both AASIST STOPA and the ResNet-34 models solely trained on STOPA produce, essentially, a chance rate. Since ASVspoof 2019 contains a larger number of training attacks, this strongly suggests that the more limited set of 3 STOPA attacks is not sufficient for capturing 'general variability' of attack-specific traits. Furthermore, using common vs. non-common trace embedding utterances has no impact on final performance.
@@
It might be useful to reflect these results on those typically obtained from speaker detection studies. Modern speaker embedding extractors, such as ECAPA-TDNN, are typically trained using thousands of speakers (classes). This leads to a discriminative 'speaker trait' space that transfers easily to classify previously unseen speakers. In contrast, we used only a handful of attacks (3 in STOPA) in training.
@@ 5 Conclusions
We introduced STOPA, a novel dataset for synthetic speech source tracing, featuring 13 distinct attacks with systematic variations in acoustic and vocoder models. Unlike existing datasets, STOPA is designed to address 'open-world' source tracing. Inspired by NIST speaker detection benchmarks, our evaluation protocols unify source tracing and unknown attack detection under a scalable framework. By prohibiting the use of other attack data for training or testing, we encourage the development of dynamic 'attack signature databases' that evolve over time, similar to antivirus systems.